OpenStudy (anonymous):

The area of a rectangle is 18 , and its diagonal is sqrt(85) . Find its dimensions and perimeter.

OpenStudy (anonymous):

let length=x width=y xy=18 x^2+y^2=sqrt(85) 2(x+y)=? x,y=? (x+y)^2=x^2+y^2+2xy =sqrt(85)+18 =27.22 (x+y)=5.2 solve this to get ur result

OpenStudy (anonymous):

im sorry im still not following

OpenStudy (anonymous):

i understand xy=18 and x^2 + y^2=sqrt(85) but where did the 2(x+y) and the rest come from?

OpenStudy (anonymous):

2(x+y) is perimeter

OpenStudy (anonymous):

x,y are dimensions

OpenStudy (anonymous):

(x+y)^2=x^2+y^2+2xy is an identity

OpenStudy (anonymous):

put values in this identity and find the unknowns